Attending: Christoph Montag, Mike Leitch, Stephen Trentalange, Fulvia Pilat,, Wolfram Fischer, Paul Sampson, Phil Pile, Yousef Makdisi, Mike Sivertz.
Paul: Usual amount of jobs. Leak in H17, Vacuum is OK, but current in the ion pumps keeps increasing. New pipes exist for replacement, but there is no consensus for the need or wisdom of replacing the pipe now. There is also the 880 Polarimeter. For now, the maintenance Day is postponed until the 27 Feb 2008.
Fulvia deferred the decision about an APEX on 12 Feb until Friday. Otherwise the first APEX is 20 Feb 2008.
Yousef needs a 2-hour access to warm up the JET for nozzle cleaning. It was last cleaned 6 days ago, so we can see how long it takes to clog up.
Christoph wants to know what the experiments need to collisions. PHENIX wants to begin with a transverse vertical store to set up their polarization equipment. They want 5-6 CNI measurements throughout the fill to calibrate their equipment. Need Angelika to set up collimation tomorrow. She will try to set up collimation from home tonight around 10:00PM.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 (leave message) |